wadecain23:

Tom, Martin, and Rocky work at a call center. Last Friday, Tom worked for 5 hours, Martin worked for 6.25 hours, and Rocky worked for 6 hours and 30 minutes. What is the total time that they worked?

JustSaiyan:

Well let's start by converting the decimals. 6.25 hours = 6 1/4. 1/4 of 60 is 15. So their work time is as follows Tom: 5 hours Martin: 6 hours and 15 minutes Rocky: 6 hours and 30 minutes 6+6+5= 17 hours 15+30= 45 minutes Total time worked: 17 hours and 45 minutes

Kayden:

the answer is 17.55 because 6.25 + 5 = 11.25 +6.30 = 17.55 6 hour and 30 min is 6.30